Prepare 50 mL of TRIS buffer at strength of 0.1 M and at a pH of 7.5 using Tris, 1 M HCl and water.

TRIS pKa: 8.3

TRIS formula weight: 121.1 g/mol
Answer #1

50 mL of TRIS buffer at strength of 0.1 M

we have,

pH = pKa + log [A]/[HA]

at a pH of 7.5 , TRIS pKa: 8.3

we get,   7.5 = 8.3 + log [A]/[HA]

log [A]/[HA] = -0.8 ==> [A]/[HA] = 0.16

since, [A]+[HA] = 0.1 M

thus, [HA] = 0.086M and [A] = 0.014 M

===>

for 50 mL of buffer ,

we need , TRIS = 0.1 mol/L * 0.05 L = 0.005 moles

molar mass tris :121.1 g/mol

thus, mass TRIS = 0.6055 g

Amount of 1M HCl need to be added : ( from HA)

=   0.086* 50 mL / 1.0 M = 4.3 mL

==> add water to make volume 50 mL.